Emergency maintenance – 10/08/2016 17:30hrs – IPStream Connect and WBC connections

Posted: Wednesday, August 10th, 2016 at 12:01 by Neil Watson

We are currently aware that some customers are experiencing a level of packet loss on IPSC and a small number of WBC nodes. As well as investigating internally, we have also engaged with 3rd party suppliers who have identified a configuration problem that will lead to the issues being experienced. In order to correct this, changes need to be made, however in doing so, this will cause a port flap on our equipment which will effectively disconnect all DSL customers on the following nodes:

All IPSC nodes and the following WBC nodes
Peterborough
Preston
Sheffield
Slough
Stepney Green
Wolverhampton

We will therefore undertake this work at approximately 17:30hrs this evening to restore full service to all customers. It is anticipated that all customer will reconnect automatically, however we would ask anyone that does not connect to power cycle their on site equipment as a first step.

We wish to apologise for the short notice of this emergency maintenance, but feel that it is necessary in the interest of the affected customers. We’d also like to apologise for any inconvenience that this may cause and wish to reassure you that we are working hard to minimise any impact.

13 Responses to “Emergency maintenance – 10/08/2016 17:30hrs – IPStream Connect and WBC connections”

  1. Neil Watson says:

    We have been advised that the 3rd party supplier is not currently in a position to undertake the changes and is working on an alternative method of resolving the issue. We will therefore cancel the maintenance at 17:30 today. Should any further work be needed we will look to roll this into the pre-planned maintenance window from 23:00 tonight. Apologies for the short notice. We will continue to work to get the issue resolved and will update further on this post.

  2. Neil Watson says:

    Work was undertaken during the maintenance window last night to correct the packet loss issue that some customers were seeing. This work did not disconnect users and did not impact users’ services negatively. Initial reports this morning indicate that the levels of packet loss are greatly diminished or are not present. We are however seeing some disruption to VoIP services which are impacting a minority of users. We are currently trying further diagnose. We apologise for the ongoing nature and would like to reassure you that the issue has high level visibility within Entanet and we are working to resolve as quickly as possible.

  3. Neil Watson says:

    Investigations are continuing and further escalations have been made with 3rd party suppliers. We are expecting a further update before the close of play today and will update as soon as we have further information.

  4. Neil Watson says:

    We have received a further update from our suppliers. They are now planning on introducing further monitoring and diagnostics to confirm the suspected cause of the VoIP issues being seen. Once confirmed, there is an anticipated root to fix which will be implemented as soon as possible. More info will be posted as soon as we have it.

  5. Neil Watson says:

    Further escalations have been made to push this issue as we’re keen to get a resolution as soon as possible. We apologise for the delays being experienced in returning full service to all customers. We will update with more concrete information once we have it.

  6. Neil Watson says:

    Investigation work has continued throughout the day and a potential problem with mis-categorised flows has been identified. Work is underway to determine that this is indeed the cause, after which plans for rectification will be implemented. We will continue to update until resolved.

  7. Neil Watson says:

    Further work has continued over the weekend and changes have been agreed with 3rd party suppliers which have been implemented. These changes appear to have resolved the issue with VoIP and our monitoring has been positive since the changes were made. Unfortunately we will only be able to confirm once we have seen the normal weekday load return. We will continue to monitor closely and would encourage partners and customers to report further issues should they be seen. Once again apologies for the incident and the time taken to resolve.

  8. Adam Heath says:

    Following the previous update, early reports from users inform us that there is still disruption being experienced on some VoIP services. We can assure you this is being handled as the highest priority and we are continuing to work with the suppliers for a resolution.

  9. Neil Watson says:

    Following on from this morning’s reports of further VoIP disruption we have, at the manufacturer’s request, applied additional configuration changes to correct the miscategorised traffic flows. This action was determined based on the traffic dumps and monitoring that took place this morning as soon as the reports of further issues were received. Initial reports following the change suggest that this has had a positive effect. We will continue to monitor and ask for partners to contact the support team should they experience further problems.

  10. Neil Watson says:

    We are receiving reports that the issue has not been resolved and that some customers are still experiencing VoIP issues. We have already passed the issue back to the 3rd party suppliers, with whom we are closely engaged – a further update is expected soon. We are also looking to make changes within the Entanet network to try to further alleviate the problem. We anticipate those changes being made shortly, after which we will re-test again.

  11. Neil Watson says:

    Since our last post we have made further changes to try to alleviate the symptoms of the issues seen. Whilst these changes will improve the quality of some calls, it is not the full solution. We are continuing to work on the root cause of the issue and will update further as soon as we are able. Once again apologies for the extended time to fix for this issue.

  12. Neil Watson says:

    Further configuration changes were made late last night. We are currently assessing the impact of those changes and will advise soon if improvement is seen. We’ll continue to work today to get to a point of resolution.

  13. Neil Watson says:

    As a result of further diagnostics and investigations we have just made further changes to the platform to alleviate the problems being experienced by some VoIP customers. We will now monitor the support calls being received for any VoIP issues that occur after this post and the interfaces of the relevant systems for any loss or drops. Based on the results of this monitoring we will then advise next steps. Additional updates will be made with the results of that monitoring.

Leave a Reply

You must be logged in to post a comment.